# A Perspective on Quantum Computing Applications in Quantum Chemistry using 25–100 Logical Qubits

### v1

@@ VI.1 Benchmark Design and Problem Class Prioritization
The need for chemically grounded benchmark problems will go beyond idealized or academic systems [27, 169]. Prioritized use cases include challenging electronic structure problems relevant to the model systems relevant to energy and materials science as described in Table 1.
Benchmarks can further be organized into classes [249] based on the evaluation metrics distinguishing moderately correlated systems tractable in the near-term from strongly correlated systems requiring large-scale fault tolerance:
Class 0 systems (weakly correlated ones) also remain important targets for verifying methods [249].
Notably, the evaluation metrics were encouraged to include quantitative agreement with experiment, convergence behavior, scaling trajectories, and chemical observables such as excitation energies, spin state gaps, and phase boundaries.
VI.2 Community Infrastructure and Interoperability
Community-driven infrastructure is considered essential for enabling reproducibility and scaling collaboration. Critical components include:
Modular, open-source codebases for hybrid quantum-classical workflows (e.g., frameworks integrating classical chemistry packages with quantum libraries [251]).
Interoperable APIs and intermediate data formats for quantum and classical backends.
Cloud-accessible simulation environments and emulators, potentially integrated with HPC ecosystems [132].
We encourage efforts to bridge the language gaps between chemists, software engineers, and hardware developers, as well as build up compiler tool chains with chemistry-aware abstractions (domain-specific languages) and automated translation between Hamiltonians and quantum circuits [251].
VI.3 Strategic Roadmap and Time-Phased Goals
Staged roadmaps have typically been discussed to guide the evolution of capabilities across the next five years, though specific timelines remain speculative [47]. Public roadmaps from hardware vendors also inform expectations [28]. Here, we emphasize realistic targets tied to hardware constraints (e.g., progress in error correction [37, 38]), software stack maturity, and community readiness, as summarized in Figure 2.
Reverse engineering of hardware requirements from target applications, such as defining chemical accuracy thresholds to determine necessary qubit counts and gate fidelities [149, 204, 252, 253], can be employed as a guiding principle. Also, defining realistic capabilities requires understanding the practical performance limits and optimization potential of near-term algorithms. For variational approaches like QML, techniques such as the stabilizer bootstrap are being developed to assess when and how these circuits might offer utilities or be efficiently optimized, informing feasibility assessments along the roadmap [223].
VI.4 Collaborative and Cross-Disciplinary Models
Progress toward these goals was linked to structured collaboration across institutions and disciplines [44]. Suggested models include:
Shared fellowships and postdoctoral exchanges across chemistry, computer science, and physics.
Hackathons and challenge problems with well-defined resource and accuracy targets.
Standardized terminology (e.g., logical vs. physical qubits, QEC-protected operations) to improve communication.
Notably, we realize that it is also important to disseminate negative results, clarify failed optimization strategies, and share classical-quantum crossover boundaries for guiding future design.
VII Conclusion and Outlook
VII.1 A Pragmatic Shift Towards Quantum Utility in Chemistry
@@
The central goal has been refined towards achieving “Quantum Utility”—demonstrating that a quantum device can outperform the best classical methods for specific, well-chosen tasks, considering metrics like speed, accuracy, efficiency, and resource costs [258]. This means prioritizing niche but meaningful problems where quantum computers are most likely to provide an impact, such as strongly correlated systems, quantum dynamics, and catalysis. Recent demonstrations, like simulating using logical qubits with error detection [259], signal tangible progress towards fault tolerance. Quantum chemistry continues to be a cornerstone application driving these advancements [142, 20, 21].
VII.2 Co-Design, Collaboration, and Building the Foundation
@@
Looking ahead, progress hinges critically on the principle of co-design. This involves deep collaboration: between algorithm development and hardware capabilities, across different stages of simulation workflows, and among diverse disciplinary communities [131, 44, 120]. Continued innovation in theory, hardware, and software is essential, particularly focusing on techniques like downfolding, low-depth subspace methods, renormalized embeddings, and novel fault-tolerant frameworks (e.g., combining fermion and qubit codes) [260]. Integrating quantum workflows with efficient classical pre-processing and optimization strategies, potentially leveraging structural properties or fault-tolerant concepts even for near-term methods [223], will also be key.
The near-term roadmap prioritizes several concrete actions:
Development of efficient classical pre-processing and optimization strategies tailored for quantum algorithms [223].
Building a robust ecosystem, including open infrastructure and community benchmarks, to ensure reproducible and reliable scientific outcomes.
Ultimately, the work ahead is inherently collaborative. Achieving production-quality quantum chemical simulations requires a shared vision, coordinated investment, and a commitment to open science [44, 120]. The current momentum indicates that the era of practical quantum chemistry enabled by quantum computers is rapidly approaching, and its foundations—built on pragmatic goals, collaborative co-design, and targeted applications—are being firmly established today.
VIII Acknowledgements
@@
This article grew out of discussions at the Workshop on Quantum Computing Applications in Quantum Chemistry, held in April 2025 in Seattle, Washington, USA. We acknowledge Dr. Nathan Baker, Mr. Brian Bilodeau, Dr. Tamas Gorbe, Dr. Hongbin Liu, and Dr. David Williams-Young from Microsoft Azure Quantum for their inputs that form fruitful discussions during the workshop. We hope this paper offers a perspective that contributes to ongoing, pressing discussions on the near-future integration of quantum computing and computational chemistry to enable new scientific advances. Finally, the authors gratefully acknowledge the agencies and programs whose support of their individual research efforts made this work possible. The authors gratefully acknowledge the agencies and programs whose support of their individual research efforts made this work possible. In particular, B.P. acknowledges the support from the Early Career Research Program by the U.S. Department of Energy, Office of Science, under Grant No. FWP 83466. M.R. is grateful to the Novo Nordisk Foundation for financial support through the Quantum for Life center in Copenhagen/Zurich, NNF20OC0059939. J.L. is supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, PQI Community Collaboration Awards, John C. Mascaro Faculty Scholar in Sustainability, NASA under award number 80NSSC25M7057, and Fluor Marine Propulsion LLC (U.S. Naval Nuclear Laboratory) under award number 140449-R08. A.L., C.L., M.Z., D.C. are supported by the ”Embedding QC into Many-body Frameworks for Strongly Correlated Molecular and Materials Systems” project, which is funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences (BES), the Division of Chemical Sciences, Geosciences, and Biosciences (under award 72689). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center, a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. Y.Z. acknowledges the support from the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of US Department of Energy (contract no. 89233218CNA000001). Finally, we thank the Quantum Algorithms and Architecture for Domain Science Initiative (QuAADS), a Laboratory Directed Research and Development (LDRD) program at PNNL, for providing the support for organizing the workshop, and in particular we are grateful to Alison Erickson for her invaluable assistance with workshop logistics and transcription.

### v2

@@ V.3 Benchmark Design and Problem Prioritization
The need for chemically grounded benchmark problems will go beyond toy or highly simplified test cases (e.g., minimal systems mainly used for proof-of-concept studies) and instead emphasize chemically realistic, experimentally relevant systems [30, 174]. Prioritized use cases include challenging electronic-structure problems tied to energy and materials science, as summarized in Table 1.
In addition to chemically motivated benchmarks, stabilizer circuits should be systematically incorporated into benchmark suites as classically verifiable checkpoints for validating compilation/execution paths and resource estimates (see Sec. III.4). Building on the benchmark framework and Class 0/1/2 definitions introduced in Sec. III.4, here we emphasize their role in guiding resource annotation and prioritization. Infrastructure benchmarks (Class 0) and challenge benchmarks (Class 1/2) serve complementary purposes within the same program: the former establish reliable execution and comparable metrics, while the latter probe computational difficulty and chemistry-relevant observables beyond current classical limits. For Class 1/2 challenge benchmarks, the protocol additionally requires classical-only convergence baselines for the classical components of the workflow (e.g., active-space size, embedding region/bath size, downfolding truncations and tensor-factorization ranks, and environment ensemble/sensitivity), together with the final quantum cost-vector (logical-qubit count, logical depth/-count, post-downfolding term counts/ norms, and shot budgets) [174, 130, 30]. These baselines justify the specific instances executed on quantum hardware and enable apples-to-apples comparisons across stacks.
Notably, the evaluation metrics are encouraged to include quantitative agreement with experiment, convergence behavior, scaling trajectories, and chemical observables such as excitation energies, spin-state gaps, and phase boundaries. Beyond these physics-based metrics, measurement and trainability should be reported explicitly on total shot budgets per observable (or per desired error bar), any measurement-reduction used (e.g., grouping, classical shadows), and trainability diagnostics (gradient norms, optimizer stability/variance scaling) under stated noise and transpilation settings. In addition, for embedding and environment-sensitive studies we recommend that benchmarks include: (i) explicit environment metadata (temperature, solvent or dielectric model, ionic composition, cofactors) and an ensemble-averaging protocol; (ii) an uncertainty budget decomposed into ensemble, solver/statistical, and model components; and (iii) emphasis on differential observables (e.g., shifts, gaps, isotope effects) with stated error bars, alongside resource annotations (see Sec. V.3). Benchmarks should be paired with provenance-rich artifacts (inputs, circuits/Hamiltonians, calibration/decoder outputs) and FAIR metadata for re-execution and auditability (see Sec. V.2).
V.4 Strategic Roadmap and Milestones
Scope: A performance-based roadmap linking algorithmic milestones with hardware/QEC, compiler/runtime, and validation metrics, emphasizing reproducibility and community-wide comparability.
What currently limits progress? A practical question for the capability-phased roadmap is what most constrains progress over the next few years. In our view, the limiting factors are dual and interdependent. On the algorithmic side, key bottlenecks include (i) efficient and portable state preparation for strongly correlated systems and real-time dynamics; (ii) the measurement overhead that scales with Hamiltonian variance and target precision (cf. Eq. (2)), motivating measurement-reduction pipelines (grouping, low-rank factorizations, classical shadows) as first-class resources; (iii) trainability and stability across geometries for variational and subspace methods (mitigated by structured/subspace ansatzes, Krylov/filtered-QPE variants, and error-aware outer loops); and (iv) robust downfolding/embedding with clear commutator truncation and self-consistent outer loops that remain stable when the fragment solver is stochastic, supported by standardized, provenance-rich benchmarks and validation protocols [13, 52, 107, 192, 193, 29, 175, 176, 177, 72, 74, 73, 53]. On the hardware side, realizing sustained computations with 25–100 logical qubits depends on attaining low logical error rates at practical code distances, adequate decoder/-factory throughput, high-fidelity entangling operations under connectivity constraints, mid-circuit measurement and classical control latencies compatible with hybrid runtimes, and reliable noise models for transpiler and QEC co-design [31, 42, 32, 43, 33, 40, 51].
Crucially, neither front can unlock the roadmap in isolation. Today, the availability and stability of error-corrected logical qubits set the primary gating constraint for end-to-end demonstrations at the 25–100 logical-qubit scale; at the same time, resource-aware algorithms (e.g., iterative/statistical/filtered QPE, measurement-efficient subspace methods, and compact effective Hamiltonians via downfolding) directly reduce hardware budgets and accelerate time-to-utility. We therefore anticipate that near-term progress will be paced by co-design: aligning algorithmic cost-vectors (logical depth/ count, post-downfolding term counts/ norms, shot budgets) with realistic QEC and runtime throughput, integrating HPC/AI for decoding, scheduling, and measurement reduction, and validating against community benchmarks with stated error bars. This interplay, rather than a single bottleneck, sets the slope of progress toward durable quantum utility.
Rather than date-specific roadmaps, we adopt performance-based milestones tied to hardware/QEC, compiler/runtime, and validation metrics (Fig. 4, see also Refs. 27, 193, 30, 226, 227). Concretely, milestones include: logical error rate per code cycle and attainable code distance (QEC progress), -factory/decoder throughput and end-to-end latency budgets (compiler/runtime), and validated observable classes at stated error bars on resource-annotated benchmarks (community readiness, see the previous section for benchmark classes) [42, 43]. In Fig. 4 and Tab. 4 we organize the roadmap by capability phases, not dates. Phase I emphasizes short-depth exemplars (e.g., variational/subspace or short-depth QPE) under selective fault-tolerance and reports code distance and logical-error-per-cycle, decoder throughput/latency, and validated observable classes for Class 0 benchmarks. Phases II and III extend these metrics to higher logical-gate depth and qubit counts, with hybrid-workflow benchmarks (Sec. IV.5) and stabilizer-based checkpoints for classically verifiable validation (Sec. III.3).
Reverse-engineering from target applications (e.g., mapping chemical-accuracy thresholds to logical-qubit counts, code distances, and gate fidelities) remains a useful design loop [159, 256, 257, 258]. Realistic capability claims should report measurable limits (trainability, measurement cost, depth at fixed code distance) under stated noise and transpilation settings. For variational circuits, stabilizer-bootstrap and classical surrogates provide classically verifiable checkpoints to quantify trainability and measurement overhead before deployment (see Sec. IV.5) [249].
Progress toward these goals depends on structured collaboration across institutions and disciplines [50]. In practice, this means cultivating shared talent pipelines and organizing multi-institutional consortia aligned with benchmark testbeds (e.g., DOE NQISRC, DARPA Quantum Benchmarking, NSF NQVL/QLCI) [259, 226, 260, 261]. Hackathons and challenge problems with specified resource and accuracy targets can accelerate reproducible comparisons and lower entry barriers. Standardized terminology (e.g., logical vs. physical qubits, QEC-protected operations) and shared documentation improve cross-domain communication. Equally crucial is the dissemination of negative results and reproducible artifacts (input decks, circuits, data schemas, calibration snapshots, decoder outputs) via living registries that pair benchmark definitions with reference implementations and provenance, enabling fair cross-platform evaluation as hardware, software, and algorithms co-evolve.
These collaborative structures provide the foundation on which the field can now chart its broader trajectory. We turn to this outlook in Sec. VI.
| Phase | QEC / hardware metrics | Compiler / runtime metrics | Validation metrics |
| I | Demonstrate target code distance (e.g., –) with logical-error-rate-per-cycle below a stated threshold. | Report end-to-end latency on critical paths and decoder throughput. | Validated observable classes for Class 0 entries (e.g., small QM9 subset, alkali hydrides, benzene) with stated error bars; include stabilizer checkpoints. |
| II | Advance to next code-distance point with lower logical error per cycle. | Provide compiler/transpiler cost model and decoder and -factory throughput at scale. | Hybrid-workflow benchmarks (e.g., Fe2S2, OLED fragments, model dynamics) with resource annotations; cross-validated vs. high-level classical baselines. |
| III | Sustained operation at higher code distance with a stable logical-error budget. | Integrated GPU–QPU orchestration targets (latency/throughput); measurement-reduction pipeline. | Chemically validated observables for Class 1 targets and/or dynamics with stated uncertainty. |
VI Conclusion and Outlook
In this concluding section, we synthesize the main insights from the preceding discussions and place them in a broader perspective. The 25–100 logical-qubit regime emerges not only as a technical milestone but also as a proving ground for collaborative strategies that will shape the trajectory of quantum chemistry in the fault-tolerant era.
VI.1 A Pragmatic Shift Towards Quantum Utility in Chemistry
@@
Building on this foundation, the central goal has been refined towards achieving “Quantum Utility,” demonstrating that a quantum device can provide reliable advantage over the best classical methods on well-chosen, domain-relevant tasks, judged by speed, accuracy, efficiency, and resource costs [263]. Rather than general-purpose superiority, this entails prioritizing niche but meaningful problems such as strongly correlated systems, quantum dynamics, and catalysis. Recent demonstrations, like simulating using logical qubits with error detection [264], signal tangible progress towards fault tolerance, with quantum chemistry continuing to be a cornerstone application [144, 20, 21]. Equally important, however, is that such demonstrations yield new scientific insights (e.g., mechanistic understanding, predictive trends, and emergent quantum behaviors) that remain elusive for classical solvers. Such insights may come, for example, from resolving the role of conical intersections in photochemistry or tracing coherence-driven pathways in enzymatic reactions, providing value beyond numerical accuracy alone and underscoring the broader impact of quantum utility on discovery.
VI.2 Co-Design, Collaboration, and Building the Foundation
@@
Looking ahead, progress hinges on co-design: tight coupling between algorithm development and hardware capabilities, coordination across the stages of hybrid simulation workflows, and sustained collaboration among chemistry, computer science, and physics communities [131, 50, 120]. Continued innovation across theory, hardware, and software remains essential, with particular emphasis on downfolding and renormalized embeddings, low-depth subspace methods, and emerging fault-tolerant frameworks (e.g., combining fermion and qubit codes) [265]. Focus should also include integrating quantum workflows with efficient classical pre-/post-processing and optimization strategies, leveraging structural diagnostics and fault-tolerant concepts (e.g., stabilizer bootstrapping) even when operating in near-term regimes [249].
In the near term, we prioritize actions that are concrete and measurable: design chemically relevant, resource-annotated benchmarks that guide development and quantify progress [174, 30]; develop efficient classical preprocessing and optimization tailored to quantum subroutines [249]; construct modular, testable hybrid workflows co-designed with hardware constraints and portable across stacks [131, 225]; and integrate AI and HPC as scalable partners for scheduling, decoding, and measurement reduction [130, 132]. These efforts should be supported by open infrastructure and community benchmarks to ensure that results are reproducible, comparable, and auditable over time.
Ultimately, the work ahead is inherently collaborative. Achieving production-quality quantum chemical simulations will require a shared vision, coordinated investment, and a commitment to open science [50, 120]. The present momentum suggests that practical quantum chemistry is within reach; by grounding progress in co-design, validated benchmarks, and targeted applications, we can build the foundation for durable utility in the 25–100 logical-qubit era and beyond. In closing, the 25–100 logical-qubit regime should be viewed as the community’s first practical window for demonstrating quantum utility in chemistry, where resource-aware algorithms, auditable benchmarks, and co-designed workflows converge to establish credible, reproducible progress beyond classical limits.
VII Acknowledgements
@@
This article grew out of discussions at the Workshop on Quantum Computing Applications in Quantum Chemistry, held in April 2025 in Seattle, Washington, USA. The authors gratefully acknowledge the agencies and programs whose support of their individual research efforts made this work possible. In particular, we acknowledge Nathan Baker, Brian Bilodeau, Tamas Gorbe, Hongbin Liu, and David Williams-Young from Microsoft Azure Quantum for their inputs that form fruitful discussions during the workshop. We hope this paper offers a perspective that contributes to ongoing, pressing discussions on the near-future integration of quantum computing and computational chemistry to enable new scientific advances. We thank the Quantum Algorithms and Architecture for Domain Science Initiative (QuAADS), a Laboratory Directed Research and Development (LDRD) program at PNNL, for providing the support for organizing the workshop, and in particular we are grateful to Alison Erickson for her invaluable assistance with workshop logistics and transcription. A.L., C.L., M.Z., D.C. are supported by the ”Embedding QC into Many-body Frameworks for Strongly Correlated Molecular and Materials Systems” project, which is funded by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences (BES), the Division of Chemical Sciences, Geosciences, and Biosciences (under award 72689). J.L. is supported in part by the University of Pittsburgh, School of Computing and Information, Department of Computer Science, Pitt Cyber, PQI Community Collaboration Awards, John C. Mascaro Faculty Scholar in Sustainability, NASA under award number 80NSSC25M7057, and Fluor Marine Propulsion LLC (U.S. Naval Nuclear Laboratory) under award number 140449-R08. B.P. acknowledges the support from the Early Career Research Program by the U.S. Department of Energy, Office of Science, under Grant No. FWP 83466. MvS is supported by the Computational Chemical Sciences program within the Office of Basic Energy Sciences, U.S. Department of Energy (DOE) under Contract No. DE-AC36-08GO28308. M.R. is grateful to the Novo Nordisk Foundation for financial support through the Quantum for Life center in Copenhagen/Zurich, NNF20OC0059939. Y.Z. acknowledges the support from the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of US Department of Energy (contract no. 89233218CNA000001).
Appendix A Extended details on qumodes and cQED gate implementations
Qumodes or quantum harmonic oscillators are known for their infinite-dimensional Hilbert space and offer a compelling resource beyond the two-level systems of conventional qubits [231]. In practice, the Fock basis of a qumode within an energy cutoff represents a discrete multi-dimensional generalization of a qubit, also known as a qudit [232].
Significant progress has been made on quantum devices consisting of qumodes coupled to a qubit, based on the cQED approach [233], where qumodes are realized as microwave cavities dispersively coupled to a superconducting transmon acting as the qubit [266]. In addition to providing more resources than qubit-only architectures, cQED devices support a diverse range of universal gate sets, including echoed conditional displacement (ECD) with qubit rotations [267], selective number-dependent arbitrary phase (SNAP) gates with displacements and beamsplitters [268], and conditional-not displacement gates [269], among others.
These gates have recently been explored in applications spanning chemistry [270, 271, 272, 273], optimization [274], and quantum machine learning [275]. Hybrid qubit–qumode gates are also difficult to mimic using shallow qubit-only circuits [236, 237], and thus offer potential advantages for problems with strong bosonic character (e.g., vibronic simulations [236]) or optimization landscapes favorable to these circuits [276].
This extended discussion provides the technical backdrop for the concise treatment of qumodes in Sec. V.1 of the main text.